Calculate the Hz0* concentration from each pH: (a) 7.42; (b) 12.14.
Answer #1

pH = 7.42 s-log[++] = -7.42 t] =-7.42 [HT] = 10 -7.42 = 3,8 x 10-8 M bi pH = 12.14 -log 10 [H+] = 12.17 logo [ht] = - 12.114